Thomas’ Calculus Early Transcendentals 12ed

Homework Assignment for Chapter 04

1. Section 4.1: problems 27, 35, 39, 59, 77, 93.

2. Section 4.2: problems 17, 18, 61, 65, 66.

3. Section 4.3: problems 7, 11, 74, 77, 80.

4. Section 4.4: problems 77, 85, 105, 111.

5. Section 4.5: problems 21, 51, 57, 65, 73, 80, 81(b), 84(c), 86.

6. Section 4.6: problems 12, 27, 48, 62, 67(a).

7. Section 4.7: problem 9. Then in problems 18, 21, 29, 30, give formula for Newton’s method, and need not find the numerical value of the solution. In problem 29, also estimate how many iterations it takes to get |x_{n} − x_{∗}| ≤ 10^{−3}. Hint for problem 18: you need to find an equation so that π is a solution. Trigonometric functions will help.

8. Section 4.8: pick at least 10 among the odd numbered ones in problems 25-69. Problems 79, 80, 84, 127, 128.